# Faraday laser pumped cesium beam clock

### v1

@@ IV Conclusion
We implement a Faraday laser pumped cesium beam clock. Applying Faraday laser to cesium beam clock has several advantages. First, Faraday laser is immune to the changes in current and temperature of the laser diode. In addition, because the Faraday laser uses atoms as frequency reference, the frequency of the Faraday laser immediately corresponds to the atomic spectra upon switching on and does not require manual tuning. Furthermore, the frequency-stabilized Faraday laser based on MTS technique has narrow linewidth and high frequency stability. Measured by optical heterodyne method, the linewidth of the frequency-stabilized Faraday laser is 2.5 kHz and the frequency stability of the frequency-stabilized Faraday laser is at 1 s due to the high SNR of the MTS signal. Thus the application of Faraday laser to cesium beam clock can improve the SNR of Ramsey fringes leadng to the optimization of the performance of cesium beam clock. When the cesium oven temperature is 110°C, the cesium beam clock achieves a SNR in 1 Hz bandwidth of , and the frequency stability of the cesium beam clock is , and decreases to at 10000 s. When the cesium oven temperature is 130°C, the cesium beam clock achieves a SNR in 1 Hz bandwidth of . This Faraday laser pumped cesium beam clock has great potential and value in the fields of timekeeping, navigation, and communication. Meanwhile, the frequency-stabilized Faraday laser in this experiment can be applied in the field of precision measurement to further improve the performance of atomic clocks, atomic gravimeters, atomic magnetometers, and other applications. In our future work, we will continue to improve the atomic utilization efficiency of the cesium beam by hexapole magnetic system[14] and two-laser pumping scheme[36] and optimize the linewidth and noise of the Faraday laser to improve the performance of the cesium beam clock.

### v2

@@ IV Conclusion
We implement a Faraday laser pumped cesium beam clock. Applying the Faraday laser to a cesium beam clock has several advantages. First, the Faraday laser is immune to the changes in current and temperature of the laser diode. In addition, because the Faraday laser uses atoms as frequency reference, the frequency of the Faraday laser immediately corresponds to the atomic spectra upon switching on and does not require manual tuning. Furthermore, the frequency-stabilized Faraday laser based on the MTS technique has narrow linewidth and high frequency stability. Measured by the optical heterodyne method, the linewidth of the frequency-stabilized Faraday laser is 2.5 kHz and the frequency stability of the frequency-stabilized Faraday laser is at 1 s due to the high SNR of the MTS signal. Thus, the application of the Faraday laser to cesium beam clock can enhance the SNR of Ramsey fringes, leading to the optimization of the cesium beam clock’s performance. When the cesium oven temperature is 110°C, the cesium beam clock achieves an SNR in 1 Hz bandwidth of , and the frequency stability of the cesium beam clock is , and decreases to at 10000 s. When the cesium oven temperature is 130°C, the cesium beam clock achieves an SNR in 1 Hz bandwidth of . This Faraday laser pumped cesium beam clock has great potential and value in the fields of timekeeping, navigation, and communication. Meanwhile, the frequency-stabilized Faraday laser in this experiment can be applied in the field of precision measurement to further improve the performance of atomic clocks, atomic gravimeters, atomic magnetometers, and other applications. In our future work, we will aim to improve the atomic utilization efficiency of the cesium beam by hexapole magnetic system[14] and two-laser pumping scheme[36] and optimize the linewidth and noise of the Faraday laser to improve the performance of the cesium beam clock.